if a system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this.
8. 3y + x = 4
x = 2y − 1

Answer by richwmiller(17219)

3y+x=4
x=4-3y
x=2y-1
4-3y=2y-1
5=5y
y=1
x=2-1
x=1
x=1 y=1